# Adversarial Robustness of Variational Autoencoders across Intersectional Subgroups

### v1

@@ 1 Introduction
Autoencoders (AEs) have emerged as a versatile method in machine learning (ML) for various tasks such as dimensionality reduction [9], representation learning [24] and data generation [25]. Application areas range from medical diagnosis [20] and medical imaging [7] to self-driving cars, which depend on precise object recognition and face recognition systems in ATMs [12] for biometric authentication. The widespread adoption of AEs, even in critical applications, naturally raises concerns about their fairness in performance across different demographic subgroups and their robustness to adversarial attacks.
@@
Despite significant progress in this area, it has become increasingly apparent that these models often learn biased representations with respect to protected attributes, such as gender or age, exhibiting discriminatory biases against minority subgroups like old women in various downstream tasks [28]. A contributing factor to this issue is the inherent biases present in the training set images [8], which often feature a high disparity in representation towards minority subgroups. This disparity not only highlights issues of fairness and inclusivity but also raises critical ethical concerns about the deployment of such technologies in diverse societies. Autoencoders are also not infallible; they are particularly susceptible to well-crafted input samples [10]. These adversarial samples, characterized by minimal, humanly imperceptible perturbations, can deceive otherwise high-performing deep learning models. This vulnerability poses significant challenges to the integrity and reliability of AEs on applications in critical domains. Variational Autoencoders (VAEs) [14, 11] have emerged as a robust alternative to their deterministic counterparts (vanilla AEs), demonstrating a higher resilience to input perturbations, especially those stemming from adversarial attacks. This robustness positions VAEs as a promising solution in the pursuit of more secure and reliable deep learning models. Nonetheless, adversaries have devised methods to exploit the resilience of VAEs by introducing minor input perturbations [29] designed to elicit substantial changes in the encoding process. Typically, these efforts aim to generate reconstructions that align closely with a specific target data point selected by the adversary, thereby compromising the model’s integrity and its representation of the original input. While the robustness of models against such adversarial strategies has been extensively studied [17, 4, 18], a gap remains in the comprehensive evaluation of model robustness across different demographics and whether there are differences in the performance.
@@
In this study, we aim to provide a comprehensive evaluation of the robustness of VAEs against non-targeted adversarial attacks across various demographic subgroups. Our initial question (Q1) examines whether there are robustness disparities among these subgroups. The second part of our research (Q2) aims to understand the factors contributing to these disparities. We evaluate the adversarial robustness of different subgroups and examine the effect of the latent space disentanglement regularization parameter on subgroup robustness inequality. Although the $\beta$-VAE with an optimal regularization parameter reduced robustness disparities, it caused many samples to resemble majority class samples. We confirmed this by analyzing variations in prediction accuracy of downstream gender and age classifiers. To understand this behavior, we explored the latent space neighborhood of samples prone to subgroup switching reconstruction, identifying embedding smoothness as a potential cause.
@@ 3.1 Problem setup
$\mathcal{SG}=\{sg=s^{1}\cap s^{2}\cap\cdots\cap s^{k}\mid\leavevmode\nobreak\ s^{i}\in\{g^{i},\overline{g^{i}}\},i=1,\cdots k\}\}$ (1)
@@ 3.2 Variational Autoencoders
$\max\limits_{\theta}\mathbf{E}_{p(z)}[p_{\theta}(x|z)]$
@@
$L(\theta,\phi)=E_{z\sim q_{\phi}(z|x)}[p_{\theta}(x|z)]-\beta\cdot D_{KL}(q_{\phi}(z|x)||p(z))$ (2)
@@ 3.3 Adversarial examples
Formally, given an input image $x$, and an auto-encoder with learned reconstruction distribution $\mathbf{E}[p_{\theta}(x|z)]$, adversarial examples are defined as a subtle perturbation/modification $\delta$, which when added to the image as $x+\delta$, maximizes the gap between expected reconstruction:
$\begin{split}&\max\limits_{\delta}\|\mathbf{E}_{z^{\prime}\sim q_{\phi}(z|x+\delta)}[p_{\theta}(x|z^{\prime})]-\mathbf{E}_{z\sim q_{\phi}(z|x)}[p_{\theta}(x|z)]\|\\ &\text{s.t.\leavevmode\nobreak\ \leavevmode\nobreak\ }\|\delta\|_{p}\leq c\end{split}$ (3)
where $c$ is a hyperparameter bounding the norm of $\delta$ to ensure a very small perturbation. This process falls under the category of non-targeted and whitebox adversarial attacks within the taxonomy of adversarial attacks [29], where the adversary has access to the trained neural network model and tries to evade the system (e.g., Dodging [22]) by learning an optimal noise using Equation 3.
@@ 4.2 Robustness evaluation for non-targeted adversarial attacks
For each instance $x\in\bigcup\limits_{sg\in\mathcal{SG}}{I}_{sg}$, we evaluate the adversarial robustness of VAE $(\mathcal{F}_{{\phi},{\theta}})$ against attacks on $x$. We use the generated optimal distortion $\delta$ from the attack generation step (Section 4.1) to get the perturbed image $x+\delta$. Then, we provide both the original image $x$ and the perturbed image $x+\delta$ as inputs to VAE and we evaluate the deviation of the adversarial output from the output on the original image as:
$\Delta_{c}=\|\mathbf{E}_{z^{\prime}\sim q_{\phi}(z|x+\delta)}[p_{\theta}(x|z^{\prime})]-\mathbf{E}_{z\sim q_{\phi}(z|x)}[p_{\theta}(x|z)]\|_{2}$ (5)
@@ 5.1.2 Training
Adam optimizer is used for training with a learning rate of 1e-4. We train three different VAE models, with $\beta=\{1,5,10\}$. Higher value of $\beta$ refers to higher emphasis of the VAE to the disentanglement factor (c.f. Section 3.2) in the latent space. However, when $\beta$ is too low or too high the model learns an entangled latent representation due to either too much or too little capacity in the latent z bottleneck  [11]. The full code for all the experiments and evaluation along with additional resources can be accessed at https://anonymous.4open.science/r/robustness_of_subgroups-A85D
@@ 5.1.3 Evaluation setting and evaluation measures
To evaluate the robustness of the subgroups, we selected 60 random samples from each subgroup. This decision was driven by resource constraints, as generating sample-specific white-box adversarial attacks incurs significant computational overhead. We evaluate the robustness of VAEs to samples of different groups by measuring the adversarial deviations according to Equation 5. For all the training we optimized adversarial perturbations according to Equation 4. Then, we report on the distribution of adversarial deviation $\Delta_{c}$ for different subgroups, according to Equation 5.
@@
In Figure  2 we show the adversarial deviation (c.f., Equation 5) vs unperturbed reconstruction loss (eq. 2) of VAE’s with different $\beta$=1,5,and 10, across different subgroup instances. Higher values of adversarial deviation correspond to lower adversarial robustness of the model against the chosen samples for robustness evaluation against a given model. Lower values of unperturbed reconstruction loss correspond to better reconstruction of the unperturbed image. We notice that the adversarial deviations of samples are notably higher for attacks against VAE with $\beta=1$ and $\beta=10$ compared to that with $\beta=5$. This observation falls inline with the literature suggesting neither low or high values of $\beta$ are good [11]. Interestingly, we also notice that samples with relatively higher reconstruction loss (0.05 - 0.06) in the unperturbed scenario (mostly populated with young men and women) does not have higher deviance ($\delta_{c}\leq 0.125$) compared to some samples (old men and women) with lower unperturbed reconstruction loss (0.02 - 0.03) with $\delta_{c}\geq 0.2$. Reconstruction loss is generally evaluated as the proxy of models quality in understanding the latent feature semantics. This finding reveals the vulnerability of such subgroups where despite low loss in reconstruction in original image, fails to understand the feature space upon adversarial perturbation.

### v2

@@ 1 Introduction
Autoencoders (AEs) have emerged as a versatile method in machine learning (ML) for various tasks such as dimensionality reduction [9], representation learning [24] and data generation [25]. The widespread adoption of AEs, even in critical applications, naturally raises concerns about their fairness in performance across different demographic subgroups and their robustness to adversarial attacks.
@@
Despite significant progress in this area, it has become increasingly apparent that these models often learn biased representations with respect to protected attributes, such as gender or age, exhibiting discriminatory biases against minority subgroups like old women in various downstream tasks [28]. A contributing factor to this issue is the inherent biases present in the training set images [8], which often feature a high disparity in representation towards minority subgroups. This disparity not only highlights issues of fairness and inclusivity but also raises critical ethical concerns about the deployment of such technologies in diverse societies. Autoencoders are also not infallible; they are particularly susceptible to well-crafted input samples [10]. The adversarial samples, characterized by minimal, humanly imperceptible perturbations, can deceive otherwise high-performing deep learning models. This vulnerability poses significant challenges to the integrity and reliability of AEs on applications in critical domains. Variational Autoencoders (VAEs) [14, 11] have emerged as a robust alternative to their deterministic counterparts (vanilla AEs), demonstrating a higher resilience to input perturbations, especially those stemming from adversarial attacks. Nonetheless, adversaries have devised methods to exploit the resilience of VAEs by introducing minor input perturbations [29] designed to elicit substantial changes in the encoding process. While the robustness of models against such adversarial strategies has been extensively studied [17, 4, 18], a gap remains in the comprehensive evaluation of model robustness across different demographics and whether there are differences in the performance.
@@
In this study, we aim to provide a comprehensive evaluation of the robustness of VAEs against non-targeted adversarial attacks across various demographic subgroups. Our initial question (Q1) examines whether there are robustness disparities among these subgroups. The second part of our research (Q2) aims to understand the factors contributing to these disparities. We evaluate the adversarial robustness of different subgroups and examine the effect of the latent space disentanglement regularization parameter on subgroup robustness inequality. Although the $\beta$-VAE with an optimal regularization parameter reduced robustness disparities, it caused many samples to resemble majority class samples. We confirmed this by analyzing variations in prediction accuracy of downstream gender and age classifiers. To understand this behavior, we explored the latent space neighborhood of samples prone to subgroup switching reconstruction, identifying lack of embedding smoothness as a potential cause.
@@ 3.1 Problem setup
$\mathcal{SG}=\{sg=s^{1}\cap s^{2}\cap\cdots\cap s^{k}\mid\leavevmode\nobreak\ s^{j}\in\{g^{j},\overline{g^{j}}\},j=1,\cdots k\}\}$ (1)
@@ 3.2 Variational Autoencoders
$\max\limits_{\theta}\mathbf{E}_{p(z)}[\log p_{\theta}(x|z)]$
@@
$L(\theta,\phi)=E_{z\sim q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-\beta\cdot D_{KL}(q_{\phi}(z|x)||p(z))$ (2)
@@ 3.3 Adversarial examples
Formally, given an input image $x$, and an auto-encoder with learned reconstruction distribution $\mathbf{E}[\log p_{\theta}(x|z)]$, adversarial examples are defined as a subtle perturbation/modification $\delta$, which when added to the image as $x+\delta$, maximizes the gap between expected reconstruction:
$\begin{split}&\max\limits_{\delta}\|\mathbf{E}_{z^{\prime}\sim q_{\phi}(z|x+\delta)}[\log p_{\theta}(x|z^{\prime})]-\mathbf{E}_{z\sim q_{\phi}(z|x)}[\log p_{\theta}(x|z)]\|\\ &\text{s.t.\leavevmode\nobreak\ \leavevmode\nobreak\ }\|\delta\|_{p}\leq c\end{split}$ (3)
where $c$ is a hyperparameter bounding the norm of $\delta$ to ensure suitable perturbation intensity. This process falls under the category of non-targeted and whitebox adversarial attacks within the taxonomy of adversarial attacks [29], where the adversary has access to the trained neural network model and tries to evade the system (e.g., Dodging [22]) by learning an optimal noise using Equation 3.
@@ 4.2 Robustness evaluation for non-targeted adversarial attacks
For each instance $x$ we evaluate the adversarial robustness of VAE $(\mathcal{F}_{{\phi},{\theta}})$ against attacks on $x$. We use the generated optimal distortion $\delta$ from the attack generation step (Section 4.1) to get the perturbed image $x+\delta$. Then, we provide both the original image $x$ and the perturbed image $x+\delta$ as inputs to VAE and we evaluate the deviation of the adversarial output from the output on the original image as:
$\Delta_{c}=\|\mathbf{E}_{z^{\prime}\sim q_{\phi}(z|x+\delta)}[\log p_{\theta}(x|z^{\prime})]-\mathbf{E}_{z\sim q_{\phi}(z|x)}[\log p_{\theta}(x|z)]\|_{2}$ (5)
@@ 5.1.2 Training
Adam optimizer is used for training with a learning rate of 1e-4. We train three different VAE models, with $\beta=\{1,5,10\}$. Higher value of $\beta$ refers to higher emphasis of the VAE to the disentanglement factor (c.f. Section 3.2) in the latent space. However, when $\beta$ is too low or too high the model learns an entangled latent representation due to either too much or too little capacity in the latent z bottleneck  [11]. The full code for all the experiments and evaluation along with additional resources can be accessed at https://github.com/ChethanKodase/robustness_of_subgroups
@@ 5.1.3 Evaluation setting and evaluation measures
To evaluate the robustness of the subgroups, we selected 60 random samples from each subgroup. This decision was driven by resource constraints, as generating sample-specific white-box adversarial attacks incurs significant computational overhead. We evaluate the robustness of VAEs to samples of different groups by measuring the adversarial deviations according to Equation 5. For all the trained models and chosen random samples we optimized adversarial perturbations according to Equation 4 by empirically setting the perturbation norm bound c for all experiments to enhance robustness comparisons. A very low c may cause no output damage in some samples, while an extreme c could cause excessive damage in all output samples, both of which hinder effective comparison. Our work aims to evaluate and compare adversarial robustness, not to demonstrate attack success. Since autoencoders are more resilient than classifiers [10], we use perceptibly intense perturbations, as also seen in [27, 10, kuzina2022alleviating], to enable robustness comparisons. Then, we report on the distribution of adversarial deviation $\Delta_{c}$ for different subgroups, according to Equation 5.
@@
In Figure  2 we show the adversarial deviation (c.f., Equation 5) vs unperturbed reconstruction loss (eq. 2) of VAE’s with different $\beta$=1,5,and 10, across different subgroup instances. Higher values of adversarial deviation correspond to lower adversarial robustness of the model against the chosen samples for robustness evaluation against a given model. Lower values of unperturbed reconstruction loss correspond to better reconstruction of the unperturbed image. We notice that the adversarial deviations of samples are notably higher for attacks against VAE with $\beta=1$ and $\beta=10$ compared to that with $\beta=5$. This observation falls inline with the literature suggesting neither low or high values of $\beta$ are good [11]. Interestingly, we also notice that samples with relatively higher reconstruction loss (0.05 - 0.06) in the unperturbed scenario (mostly populated with young men and women) does not have higher deviance ($\Delta_{c}\leq 0.125$) compared to some samples (old men and women) with lower unperturbed reconstruction loss (0.02 - 0.03) with $\Delta_{c}\geq 0.2$. Reconstruction loss is generally evaluated as the proxy of models quality in understanding the latent feature semantics. This finding reveals the vulnerability of certain subgroups where, despite the low reconstruction loss of the original image, the model fails to comprehend the feature space and resist adversarial perturbations against samples belonging to these subgroups.
@@ 7 Conclusion
8 Acknowledgements
This research work received fund from the European Union under the Horizon Europe MAMMOth project, Grant Agreement ID: 101070285, and also supported by the EU Horizon Europe project STELAR, Grant Agreement ID: 101070122